OpenStudy (anonymous):

If cos2theta = 1/3 and 0 ≤ theta ≤ pi/2, find costheta.

OpenStudy (anonymous):

\[\cos^2(\theta) or \cos(2\theta)\]?

OpenStudy (anonymous):

Cos(2theta) = 2 cos^2 theta -1

OpenStudy (anonymous):

cos(2θ)

OpenStudy (anonymous):

use a calculator enter: \[\cos(\frac{\cos^{-1}(1/3)}{2})\]

OpenStudy (anonymous):

.816 radians

OpenStudy (anonymous):

If u are expected to provide an exact value use the equation I gave above.

OpenStudy (anonymous):

yeah i needed the exact value, thanks so you do 2cos^2theta - 1 = 1/3 and solve for costheta, got it